# Token Preference Optimization with Self-Calibrated Visual-Anchored Rewards for Hallucination Mitigation

### v1

@@ 4.3 Ablation Studies
To further demonstrate how our method helps improve model performance by reducing hallucinations, we conduct ablation experiments for the noisy steps, the introduced parameter $a$, and the self-calibrated process for rewards. These experiments are performed on LLaVA-1.5-7B consistently with the main experiments, including hallucination and general evaluation benchmarks.
Noise Steps : We ablate on the noise steps in Figure 3 (a). As shown, the optimal performance is achieved at the step of 500. This medium corruption enables the model to grasp the general outline of the image while missing the detailed contents, which is prone to generate hallucinations of the visual-anchored tokens.
Parameter $a$ : We conduct experiments by varying the parameter $a$ introduced in Equation (6) with the results shown in Figure 3 (b). By setting $a=[0,0.5,1]$, we observed consistently good performance across all configurations. This suggests that effective performance is achieved as long as the reward mechanism successfully highlights token differences and identifies visually anchored tokens. Notably, the best overall results are obtained with $a=0.5$, validating our proposed method and hypothesis. This indicates that when the visual-anchored score $s=0$, setting $c=1$, not introducing additional reward signals can yield better outcomes.
Visual-Anchored Rewards : Furthermore, Table 3 demonstrates that TPO can enhance model performance when rewards are assigned separately to positive and negative samples, achieving results comparable to those obtained by rewarding both simultaneously. However, by providing opposite rewards to positive and negative samples, where rewards are negatively correlated with the visual relevance of positive samples and positively correlated with that of negative samples, TPO’s performance significantly deteriorates. In some metrics, this approach yields even poorer results than the original LLaVA-1.5 model. This further underscores the validity of the designation of visual-anchored rewards.
By integrating these selection methods, we achieve superior overall performance, thereby highlighting the value and efficacy of our approach. This integration underscores both the simplicity and effectiveness of our method in achieving robust results.
@@ 4.4 Analysis
Attentions : To further substantiate the effectiveness of our method in enhancing visual alignment, we visualize the relevance between response and image tokens, both before and after training with TPO. In this statistic, the relevance is quantified using the sum of attention weights. As depicted in Figure 4, post-TPO training, the overall image attention weights for response tokens significantly increase, especially for visual-anchored tokens (e.g., table, cord). This underscores our method’s efficacy in improving the model’s capacity to integrate visual information during response generation, thus mitigating the hallucination issue.
Self-Calibration : To illustrate that our method enables the model to progressively enhance its focus on visual information through continuous self-calibration during training, we present the evolution of scores for positive and negative samples, as calculated by Equation (6), across various training steps. With $a=0.5$, it follows that $c_{y_{i}}\in(0.5,1.5)$. As shown in Figure 5, the scores for positive samples gradually approach their maximum values, while those for negative samples approach their minimum values, indicating convergence. This trend illustrates the self-calibrating effect of our method, which ultimately enhances the model’s ability to focus on visual information.
@@ 5 Conclusion
In this study, we propose TPO, a novel token preference optimization framework, aimed at mitigating hallucinations in LVLMs. TPO incorporates a self-calibrated visual-anchored reward mechanism that automatically identifies "vision-anchored tokens" and adaptively assigns appropriate rewards to them. By adding noise to the visual input and capturing changes in the generation probability of each token, TPO computes a score indicating each token’s relevance to visual information. Subsequently, a self-calibration process adjusts these scores to ensure that: 1) Rewards for positive examples exceed those for negative examples. 2) The final rewards enhance the focus on visual information for all of "vision-anchored tokens" in preference data. Based on the self-calibrated visual-anchored reward, TPO can perform more efficient token-level preference alignment optimization for LVLMs. Experimental results have proved that TPO not only alleviates the hallucination problem but also strengthens the model’s attention to visual input when generating responses.
@@ 7 Ethic Statement
Appendix A Setup Details
@@
In our experiments, we trained the LLaVA-v1.5 model. For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-8 on the 7B version and trained for 4 epochs. We set the maximum learning rate to 2e-7 on the 7B version and trained for 4 epochs. The RLHF-V training was set according to the paper Yu et al. (2024b). All parts requiring GPT-4 evaluation use the GPT-4-0613 version, and the MM-Vet testing is conducted on the official evaluation website.
@@
Our experiments were all conducted on a server equipped with 8 GPUs; in specific cases (such as the 13B model), we utilized 32 GPUs. For the hyperparameter setting, all hyperparameters are consistent with those of our main experiment. Moreover, the level of diffusion noise in our model is represented by a formula $\xi=Sigmoid(l_{t})\times(0.5\times 10^{-2}-10^{-5})+10^{-5}$, where $l_{t}$ is a list of 1,000 numbers taken at equal intervals over the interval $[-6,6]$.
@@
The cases in Figure 1 and Figure 4 come from benchmarks Sun et al. (2023), while the cases in Figure 2 come from the RLHFV training set Yu et al. (2024b).
Appendix B Experimental Results
In this section, we will provide a detailed presentation of further ablation studies on the noise steps and $a$ in Eq 6, as shown in Table 4 and Table 5. The optimal performance for noise steps is attained at the 500-steps. This level of medium corruption allows the model to capture the general outline of the image while neglecting finer details, which can lead to the generation of hallucinations in the visual-anchored tokens. and The Optimal performance for $a$ is $0.5$, when the visual-anchored score $s=0$, we get $c=1$, not introducing additional reward signals can yield better outcomes.

### v3

@@ 4.4 Ablation Studies
Visual-Anchored Rewards
Table 4 demonstrates that TPO can enhance model performance when rewards are assigned separately to positive and negative samples, achieving results comparable to those obtained by rewarding both simultaneously. However, by providing opposite rewards to positive and negative samples, where rewards are negatively correlated with the visual relevance of positive samples and positively correlated with that of negative samples, TPO’s performance significantly deteriorates. In some metrics, this approach yields even poorer results than the original LLaVA-1.5 model. This further underscores the validity of the designation of visual-anchored rewards.
Hyperparameters
To optimize the hyperparameters in TPO, we perform comparative experiments on the noise steps (Section 3.2) and parameter $a$ in Equation 5. As the Figure 5 shown, the model performs best with 500 noise steps. Testing $a$ across the range $a=[0,0.5,1]$, we find that $a$=0.5 produces the best outcome. This supports our hypothesis that setting $s=0$ and $c=1$, without adding extra reward signals, leads to superior performance [^2] The more detail results and discussions are provided in Appendix..
@@ 4.5 Analysis
Visual-Anchored Rewards
As Figure 1 shown, the proposed visual-anchored rewards can reflect the degree to which a token depends on visual information. To further prove this statement, we construct the analysis experiment on the MMhal dataset as shown in Table 5. Intuitively, nouns and adjectives in responses are thought to most associate the content of an image. Therefore, we first perform part-of-speech (POS) tagging on the model responses and count the average number of noun/adjective tokens and other types of tokens. Specifically, in the ground-truth responses, 39.6% of the tokens are nouns or adjectives. In the responses from LLAVA-1.5-7B, the proportion of noun and adjective tokens remains nearly constant at 39.2%, both before and after TPO.
Afterwards, we count the average score from Equation 4 of noun/adjective tokens and other types of tokens. The results show that noun and adjective tokens have significantly higher scores than other types, indicating higher relevance to images. After applying TPO, these scores of all the tokens increased notably. The results supports our conclusions: 1) The visual-anchored rewards reflects token-image relevance. 2) TPO enhances the alignment of generated tokens with image content.
Attentions
To further validate TPO’s effectiveness in enhancing visual alignment, we measure the relevance using the sum of attention weights between responses and images. On the MMHal dataset, the overall image attention weights for LLaVA-1.5-7B increased from 0.14 before TPO training to 0.17 afterward. Additionally, Figure 3 visualizes the cases, showing a significant increase in image attention weights for response tokens, especially for visual-anchored tokens (e.g., table, cord). This highlights our method’s success in improving the model’s integration of visual information, thus reducing hallucinations.
Self-Calibration
To illustrate that our method enables the model to progressively enhance its focus on visual information through continuous self-calibration during training, we present the evolution of scores for positive and negative samples, as calculated by Equation (5), across various training steps. With $a=0.5$, it follows that $c_{y_{i}}\in(0.5,1.5)$. As shown in Figure 4, the scores for positive samples gradually approach their maximum values, while those for negative samples approach their minimum values, indicating convergence. This trend illustrates the self-calibrating effect of our method, which ultimately enhances the model’s ability to focus on visual information.
@@ 5 Conclusion
In this study, we propose a novel pereference alignment method, TPO, to mitigation hallucinations in LVLMs. TPO incorporates a self-calibrated visual-anchored reward mechanism that automatically identifies "vision-anchored tokens" and adaptively assigns appropriate rewards to them. By adding noise to the visual input and capturing changes in the generation probability of each token, TPO computes a score indicating each token’s relevance to visual information. Subsequently, a self-calibration process adjusts these scores to ensure that: 1) Rewards for positive examples exceed those for negative examples. 2) The final rewards enhance the focus on visual information for all of "vision-anchored tokens" in preference data. Based on the self-calibrated visual-anchored reward, TPO can perform more efficient token-level preference alignment optimization for LVLMs. Experimental results have proved that TPO not only alleviates the hallucination problem but also strengthens the model’s attention to visual input when generating responses.
@@ 7 Ethic Statement
Appendix A Implement Details
A.1 Setup
@@
In our experiments, we trained the LLaVA-v1.5 model. For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-8 on the 7B version and trained for 4 epochs. We set the maximum learning rate to 2e-7 on the 13B version and trained for 4 epochs. The RLHF-V training was set according to the paper Yu et al. (2024b). All parts requiring GPT-4 evaluation use the GPT-4-0613 8K version, and the MM-Vet testing is conducted on the official evaluation website.
We also trained the Qwen2-vl model. For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-9 and trained for 4 epochs.
For a fair comparison, we set the seed to 42 during training and greedy decoding was used during inference.
@@
Our experiments were all conducted on a server equipped with 8 Nivdia A100 GPUs; in specific cases (such as the 13B model), we utilized 32 Nivdia A100 GPUs. For the hyperparameter settings, all hyperparameters are consistent with those of our main experiment. Moreover, the level of diffusion noise in our model is represented by a formula $\xi=Sigmoid(l_{t})\times(0.5\times 10^{-2}-10^{-5})+10^{-5}$, where $l_{t}$ is a list of 1,000 numbers taken at equal intervals over the interval $[-6,6]$, and $\epsilon\in N(0,1)$.
@@
The cases in Figure 1 and Figure 3 come from benchmarks Sun et al. (2023), while the cases in Figure 2 come from the RLHF-V training set Yu et al. (2024b).
A.2 Benchmarks
The three hallucination benchmarks: (1) AMBER : a multi-dimensional hallucination benchmark with more than 15K samples, including discriminative and description tasks. (2) MMHal-Bench : it measures the hallucination rate and informativeness of responses. (3) HallusionBench : it evaluates visual illusions and knowledge hallucinations through systematically structured discriminative tasks.
The four general benchmarks: (1) SEED Bench : a benchmark for LVLMs on generative comprehension. (2) MMBench: a comprehensive benchmark designed to evaluate the capabilities across various tasks and modalities. (3) LLaVA Bench: a benchmark for evaluating multi-modal conversation, detailed description, and complex reasoning. (4) MM-Vet: a benchmark to assess integrated capabilities.
A.3 Training Efficiency
In TPO, generating corrupted images at each step incurs almost no time cost, as it is done during the initial data preparation. The main time consumption comes from calculating logits $p_{log}(y_{i}|x,v_{c},y_{<i})$ for the noisy images.
We have also conducted a careful analysis of the time consumption for LLava-1.5-7B under the settings in Section A.1, the training durations for DPO and TPO were 1 hour 24 minutes and 1 hour 57 minutes, respectively, indicating about a 40% increase in time. Nevertheless, all training methods aimed at eliminating hallucinations inevitably incur additional time costs, compared to other methods requiring fine-grained annotations, our self-calibrated approach with 40% time increase proves to be sufficiently efficient.
It has also shown superior outcomes on 5K training data training to CSR training on 13K data and POVID training on 17K data. This highlights the efficacy of our method in guiding the model to pay more attention to image details and in reducing hallucinations. We promise we will further elaborate on our efficiency in detail in the final version.
Appendix B Ablation Analysis
Noise Step
We ablate on the noise steps in Figure 5 (a). As shown, the optimal performance is achieved at the step of 500. This medium corruption enables the model to grasp the general outline of the image while missing the detailed contents, which is prone to generate hallucinations of the visual-anchored tokens.
The Figure also shows when step=0, TPO still effective and significantly better than DPO. This confusion is a code-implementation issue. In implementation as shown in Listing 1, we first convert the image into a tensor, add noise, and then convert it back into an image. This encode-decode process introduces some losses. Our method of setting the noise step to 0 serves as an ablation experiment to test the impact of this loss on our method, and it allows our experiment to more comprehensively demonstrate the advantages of TPO. The following portion of code may help you better understand our encode-decode process for adding noise. We will also open source all the code once the paper is accepted.
Parameter $\bf a$
We conduct experiments by varying the parameter $a$ introduced in Equation (5) with the results shown in Figure 5 (b). By setting $a=[0,0.5,1]$, we observed consistently good performance across all configurations. This suggests that effective performance is achieved as long as the reward mechanism successfully highlights token differences and identifies visually anchored tokens. Notably, the best overall results are obtained with $a=0.5$, validating our proposed method and hypothesis. This indicates that when the visual-anchored score $s=0$, setting $c=1$, not introducing additional reward signals can yield better outcomes.
pil_to_tensor = transforms.ToTensor()
tensor_to_pil = transforms.ToPILImage()
image = Image.open(default_image_path).convert("RGB")
image_tensor = pil_to_tensor(image)
image_noisy = add_diffusion_noise(image_tensor, 500)
image_noisy = tensor_to_pil(image_noisy)
Listing 1: Example Python Code for Noise Addition
Appendix C Comparison with Decoding-based Methods
In addition to preference alignment, many researches have improved the decoding strategy of the model to eliminate hallucinations. In this section, we selected some typical works on decoding strategies, such as VCD Leng et al. (2024), OPERA Huang et al. (2024) and LURE Zhou et al. (2023), and compared the results with them. The results show that TPO achieves more significant hallucination reduction.
Preference alignment and decoding strategies are two important and parallel categories of methods for hallucination mitigation. We believe that training with preference alignment offers several advantages: 1) Direct Optimization of Output Preferences: This approach directly optimizes the model’s output to align with desired preferences without requiring changes to the decoding strategy. 2) Higher Inference Efficiency: Preference alignment typically results in more efficient inference, as it does not introduce additional complexity during the decoding process.
One key advantage of decoding methods is that they do not require retraining the model, making them highly efficient for deployment. However, this does not preclude the benefits of preference alignment. In fact, we believe combining these two approaches can yield even better results.
Appendix D Results on Object Hallucination
In the AMBER benchmark, there is a subset for evaluating object hallucinations in image description tasks. Since this paper focuses on visual question answering, this part of the experiment is included in this section. To assess the proportion of object hallucinations in image descriptions, AMBER uses Chair as the metric.
The results are shown in Figure 6. Note that ’Chair’ represents the hallucination ratio, where a smaller value indicates better model performance. To more clearly illustrate the comparison between methods in the figure, we use $10-chair$ as the indicator. The results show that TPO can not only mitigate the hallucination in visual question answering, but also eliminate the object hallucination in image descriptions to a certain extent.
Method AMBER MMHal
[…]
VCD 71.8 74.9 2.12 54.20
LURE 73.5 77.7 1.64 60.40
OPERA 75.2 78.3 2.15 54.20
LLaVA-1.5-7B 71.7 74.3 2.01 61.46
+DPO 77.5 82.1 2.14 58.33
+TPO 79.3 85 2.47 51.04
Table 8: Comparison of Results
Figure 6: Chair Performance Comparison.
Method AMBER MMHal HallusionBench
[…]
LLaVA-1.5-7B 71.7 74.3 2.01 61.5 42.6 41.2 47.2
+TPO (white) 78.0 82.7 2.26 55.2 44.2 45.4 49.3
+TPO 79.3 85.0 2.5 51.0 41.8 48.4 50.2
Table 9: Comparison of different noise adding method. “white" indicates that blank images are used in place of noisy images.
Appendix E Comparison of Different Noise Adding Methods.
To evaluate the impact of different methods of adding noise to images on our approach, we test a scheme where noise images were replaced with white images under the same experimental conditions. The results, shown in Table 9, demonstrate the superior performance of our method. We believe that the noise addition method used in our paper can control noise levels to create images that are more likely to induce hallucinations in the model, thereby achieving better results.